# Robust Semiparametric Graphical Models with Skew-Elliptical Distributions

### v2

@@ 2.3 The meta-elliptical distribution
The meta-elliptical distribution introduced in Han and Liu, (2014) relaxes the assumption of elliptically distributed data, in a similar way as the nonparanormal of Liu et al., (2009) does for Gaussian distributions. The meta-elliptical family is a strict extension of the nonparanormal distribution.
@@
(Han and Liu, (2014), Fang, (2018)) A random vector $\boldsymbol{X}$ has an elliptical distribution, $\boldsymbol{X}\sim\mathcal{E}C_{p}(\mu,\Sigma,\xi)$, if $\boldsymbol{X}$ has the following stochastic representation: $\boldsymbol{X}\overset{d}{=}\mu+\xi AU$. Where: $\mu\in\mathbb{R}^{p}$, $q:=rank(A)$, AA is a $p\times q$ matrix such that $\Sigma=AA^{\top}$, $\xi\geq 0$ is a random variable independent of $U$, $U\in\mathbb{S}^{q-1}$, which is uniformly distributed of the unit sphere in $\mathbb{R}^{q}$.
@@
(Han and Liu, (2014)) A continuous random vector $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$ has a meta-elliptical distribution, denoted by $\boldsymbol{X}\sim M\mathcal{E}_{p}(\Sigma^{0},\xi;f)$, if there exist a set of univariate monotone functions $f_{1},\ldots,f_{p}$ such that $(f_{1}(X_{1}),\ldots,f_{p}(X_{p}))^{\top}\sim\mathcal{EC}_{p}(0,\Sigma^{0},\xi)$, where $\Sigma^{0}\in\mathcal{R}_{p}$ is the latent correlation matrix.
For the distributions that belong to the meta-elliptical family, Kendall’s tau statistic in Lemma 2.1 can be exploited to estimate the unknown correlation matrix. Nevertheless, the meaning of a missing edge is different for the nonparanormal case. Indeed, when $\boldsymbol{X}\sim\operatorname{NPN}(\Sigma^{0},f)$ a missing edge implies conditional independence. On the other hand, when $\boldsymbol{X}\sim M\mathcal{E}(\Sigma^{0},\xi,f)$ a missing edge means uncorrelation.
@@
We recall the family of $\operatorname{SU\mathcal{E}}$ distributions described in Arellano-Valle and Genton, 2010b . If a random vector $\boldsymbol{X}$ has a $\operatorname{SU\mathcal{E}}$ distribution, we write $\boldsymbol{X}\sim\operatorname{SU\mathcal{E}}_{p,p}(\mu,\Sigma,\Lambda,h^{(2p)},0,\Gamma)$, where $h^{(2p)}$ is the density generator. We focus on two specifications of the family, which are: the closed skew-normal (CSN) distribution of Domìnguez-Molina et al., (2003) [^1] We build on the closed skew-normal (CSN) distribution, which represents a reparametrization of the unified skew-normal (SUN) distribution employed by Arellano-Valle and Genton, 2010b . and the unified skew-$t$ (SUT) distribution (Arellano-Valle and Genton, 2010a ). Table 1 shows how the $\operatorname{SU\mathcal{E}}$ distribution reduces into CSN and SUT distributions. Formally, we define the $\operatorname{SU\mathcal{E}}$ distribtuion.
@@
(Arellano-Valle and Genton, 2010b, ) A continuous p-dimensional random vector $\boldsymbol{X}$ has a multivariate unified skew-elliptical $\operatorname{(SU\mathcal{E})}$ distribution, denoted by $\boldsymbol{X}\sim SU\mathcal{E}_{p,p}(\mu,\Sigma,\Lambda,h^{(2p)},0,\Gamma)$, if its density function at $x\in\mathbb{R}^{p}$ is
$f_{X}(x)=2^{p}f_{p}\left(x;\mu,\Sigma,h^{(p)}\right)F_{p}\left(\Lambda z;\Gamma,h_{Q(z)}^{(p)}\right),$ (4)
where $\mu$ is the location parameter, $\Sigma$ is a $p\times p$ dispersion matrix, $h^{(p)}$ is the density generator, $F_{r}\left(x;\Sigma,h^{(r)}\right)$ the $r$-dimensional centered elliptical cumulative distribution function with $r\times r$ dispersion matrix $\Sigma$ and density generator $h^{(r)}$. Additionally, $z=\sigma^{-1}(x-\mu)$, $\Lambda$ is a $p\times p$ real matrix controlling shape, $Q(z)=z^{\top}(\Sigma^{0})^{-1}z$, $\Gamma$ is a $p\times p$ positive definite symmetric matrix, $f_{p}\left(x;\mu,\Sigma,h^{(p)}\right)=$ $|\Sigma|^{-1/2}h^{(p)}(Q(z))$ denotes the density function of an elliptically contoured distribution with location $\mu\in\mathbb{R}^{p}$. In conclusion, $h_{Q(z)}^{(p)}(u)=h^{(2p)}(u+Q(z))/h^{(p)}(Q(z))$.
@@
(Arellano-Valle and Genton, 2010b ) Given $\boldsymbol{X}\sim\operatorname{SU\mathcal{E}}_{p,p}(\mu,\Sigma,\Lambda,h^{(2p)},0,\Gamma)$, we can write $\boldsymbol{X}$ as a mixture of a CSN distribution: $\boldsymbol{X}=\mu+\sigma V_{0}^{-1/2}\boldsymbol{Z}_{0}$, where $V_{0}$ is the cumulative distribution function of a uni-dimensional random variable, called mixture variable, such that: $V_{0}(0)=0$, and $V_{0}$ is independent of $\boldsymbol{Z}_{0}\sim\operatorname{CSN}_{p,p}(0,\Sigma^{0},\Lambda,0,\Gamma)$.
@@
(Arellano-Valle and Genton, 2010b, ) If  $\text{ $V_{0}$}\sim Gamma(\nu/2,\nu/2)$, then
$\boldsymbol{X}\sim\operatorname{SU}\mathcal{E}_{p,p}(\mu,\Sigma,\Lambda,h^{(2p)},0,\Gamma)$ reduces to $\operatorname{SUT}$ distribution: $\boldsymbol{X}\sim\operatorname{SUT}_{p,p}(\mu,\Sigma,\Lambda,\nu,0,\Gamma)$.
We construct an undirected graph on the closed skew-normal distribution, following Zareifard et al., (2016).
@@
Consider an undirected decomposable graph $G=(V,E)$, a random vector $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$ is called a skew decomposable graphical model concerning the graph G with mean $\mu$, the precision matrix $\Omega$ and the skewness parameter $\alpha\in\mathbb{R}^{p}$ if its density is $\operatorname{CSN}_{p,p}(0,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$ or equivalently
@@
where $L$ and $D_{\kappa}$ are the matrix resulting from the modified Cholesky of $\Omega=LD_{\kappa}L^{\top}$, and zero in $L$ matrix determines the conditional independence relations between the $X_{1},\ldots,X_{p}$ variables, in other words: $L_{ij}=0\Leftrightarrow\Omega_{ij}=0$. Further, when the matrix $D_{\alpha}=0$, the random vector $\boldsymbol{X}\sim\operatorname{CSN}_{p,p}(0,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$ reduces to $\boldsymbol{X}\sim\operatorname{N}_{p}(0,\Omega^{-1})$.
The graphical model for the unified skew-elliptical distribution can be constructed by utilizing the sparsity structure of the closed skew-normal distribution. We rely on a semiparametric approach that considers the class of unified skew-elliptical distributions to extend the SKEPTIC estimator of Liu et al., 2012a proposing a robust estimator of graphical models that accounts for skewness and heavy tails.
@@ 3 The elliptical skew-(S)KEPTIC estimators
(Meta skew-elliptical distribution) A continuous random vector $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$ has meta skew-elliptical distribution, denoted by $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Sigma^{0},\Lambda,h^{(2p)},0,\Gamma;f)$, if there exists a set of monotone univariate functions $f_{1},\ldots,f_{p}$ such that
$(f_{1}(X_{1}),\ldots,f_{p}(X_{p}))^{\top}\sim\operatorname{SU\mathcal{E}}_{p,p}(0,\Sigma^{0},\Lambda,h^{(2p)},0,\Gamma),$ (6)
@@
where: $\boldsymbol{U}\sim HN_{p}(0,I_{p})$ and $\boldsymbol{V}\sim N_{p}(0,I_{p})$, where $HN_{p}(0,I_{p})$ is a $p$-dimensional half-normal distribution, and $N_{p}(0,I_{p})$ is a $p$-dimensional normal distribution. The distributions of $\boldsymbol{U}$ and $\boldsymbol{V}$ are independent. The first two moments of $\boldsymbol{X}$ are:
${E}(\boldsymbol{X})$ $=L^{\top}D_{\kappa}^{-\frac{1}{2}}D,$
$\text{cov}(\boldsymbol{X})$ $=\left(L^{\top}D_{\kappa}^{\frac{1}{2}}(I-D^{2})D_{\kappa}^{\frac{1}{2}}L\right)^{-1},$ (8)
where $D=\sqrt{\frac{2}{\pi}}D_{\alpha}(I_{p}+D^{2}_{\alpha})^{-\frac{1}{2}}$.
@@
Assume that $\boldsymbol{X}$ follows a meta skew-elliptical distribution: $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Omega^{-1},D_{\alpha}L,h^{(2p)},0,D_{\kappa};f)$, such that $\boldsymbol{Y}=f(\boldsymbol{X})$ denotes the unified skew-elliptical distribution $\boldsymbol{Y}\sim\operatorname{SU\mathcal{E}}_{p,p}(0,\Omega^{-1},D_{\alpha}L,h^{(2p)},0,D_{\kappa})$. The elliptical skew-KEPTIC estimator is:
${S}^{\tau,\alpha}_{ij}=\begin{cases}2\operatorname{sin}\left(\frac{\pi}{2}A_{{\tau}_{ij}}\right)&i\neq j\\ 1&i=j\end{cases}.$
As discussed in Hult and Lindskog, (2002), Spearman’s rho is not invariant in the class of elliptical distributions, and its relation with the Pearson correlation coefficient in Lemma 2.1 holds if and only if the distribution is nonparanormal. The proof of the Theorem is in Appendix A.
@@
(elliptical skew-SKEPTIC estimator) Assume that $\boldsymbol{X}$ follows a meta skew-elliptical distribution: $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Omega^{-1},D_{\alpha}L,h^{(2p)},0,D_{\kappa};f)$, such that $\boldsymbol{Y}=f(\boldsymbol{X})$ denotes the unified skew-elliptical distribution. When $\boldsymbol{Y}\sim\operatorname{SU\mathcal{E}}_{p,p}(0,\Omega^{-1},D_{\alpha}L,h^{(2p)},0,D_{\kappa})$ reduces to the CSN distribution in (5), the elliptical skew-SKEPTIC estimator is:
${S}^{\rho,\alpha}_{ij}=\begin{cases}4\sin\left(\frac{\pi}{6}A_{{\rho}_{ij}}\right)&\text{if }i\neq j\\ 1&\text{if }i=j\end{cases},\quad\quad{S}^{\tau,\alpha}_{ij}=\begin{cases}2\operatorname{sin}\left(\frac{\pi}{2}A_{{\tau}_{ij}}\right)&i\neq j\\ 1&i=j\end{cases}.$
We generally refer to the estimators in Theorem 3.4 and Lemma 3.5 as elliptical skew-(S)KEPTIC estimators. To obtain the correlation matrix using the elliptical skew-(S)KEPTIC estimators, we need to estimate the parameters of the CSN distribution. The estimation procedure is based on the EM algorithm of Abe et al., (2021) [^2] The estimated parameters come from the skew-normal distribution of Azzalini and Capitanio, (1999), with an adequate parametrization explained in Arellano-Valle and Azzalini, (2006) we conform the parameters to a CSN distribution..

### v3

@@ 2.3 The meta-elliptical distribution
The meta-elliptical distribution introduced in han2014scale relaxes the assumption of elliptically distributed data, in a similar way as the nonparanormal of liu2009nonparanormal does for Gaussian distributions. The meta-elliptical family is a strict extension of the nonparanormal distribution.
@@
(han2014scale, fang2018symmetric) A random vector $\boldsymbol{X}$ has an elliptical distribution, $\boldsymbol{X}\sim\mathcal{E}C_{p}(\mu,\Sigma,\xi)$, if $\boldsymbol{X}$ has the following stochastic representation: $\boldsymbol{X}\overset{d}{=}\mu+\xi AU$. Where: $\mu\in\mathbb{R}^{p}$, $q:=rank(A)$, AA is a $p\times q$ matrix such that $\Sigma=AA^{\top}$, $\xi\geq 0$ is a random variable independent of $U$, $U\in\mathbb{S}^{q-1}$, which is uniformly distributed of the unit sphere in $\mathbb{R}^{q}$.
If $\xi$ has a density, then $\boldsymbol{X}$ has a density of the form:
$f_{p}\left(x;\mu,\Sigma,h^{(p)}\right)=|\Sigma|^{-1/2}h^{(p)}((x-\mu)^{\top}\Sigma^{-1}(x-\mu)),$ (4)
where $h^{(p)}$ is a scale function uniquely determined by the distribution of $\xi$, and it is called density generator.
@@
(han2014scale) A continuous random vector $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$ has a meta-elliptical distribution, denoted by $\boldsymbol{X}\sim M\mathcal{E}_{p}(\Sigma^{0},\xi;f)$, if there exist a set of univariate monotone functions $f_{1},\ldots,f_{p}$ such that $(f_{1}(X_{1}),\ldots,f_{p}(X_{p}))^{\top}\sim\mathcal{EC}_{p}(0,\Sigma^{0},\xi)$, where $\Sigma^{0}\in\mathcal{R}_{p}$ is the latent correlation matrix.
For the distributions that belong to the meta-elliptical family, Kendall’s tau statistic in Lemma 2.1 can be exploited to estimate the unknown correlation matrix. Nevertheless, the meaning of a missing edge is different for the nonparanormal case. Indeed, when $\boldsymbol{X}\sim\operatorname{NPN}(\Sigma^{0},f)$ a missing edge implies conditional independence. On the other hand, when $\boldsymbol{X}\sim M\mathcal{E}(\Sigma^{0},\xi,f)$, if $\Omega_{ij}=0$, then $X_{i}$ and $X_{j}$ are conditionally uncorrelated (rossell2021dependence).
@@
We recall the family of $\operatorname{SU\mathcal{E}}$ distributions described in arellano2010skew-e. If a random vector $\boldsymbol{X}$ has a $\operatorname{SU\mathcal{E}}$ distribution, we write $\boldsymbol{X}\sim\operatorname{SU\mathcal{E}}_{p,p}(\mu,\Sigma,\Lambda,h^{(p)},0,\Gamma)$. We focus on two specifications of the family, which are: the closed skew-normal (CSN) distribution of dominguez2003multivariate [^1] We build on the closed skew-normal (CSN) distribution, which represents a reparametrization of the unified skew-normal (SUN) distribution employed by arellano2010skew-e. and the unified skew-$t$ (SUT) distribution (arellano2010skew-t). Table 1 shows how the $\operatorname{SU\mathcal{E}}$ distribution reduces into CSN and SUT distributions. Formally, we define the $\operatorname{SU\mathcal{E}}$ distribtuion.
@@
(arellano2010skew-e) A continuous p-dimensional random vector $\boldsymbol{X}$ has a multivariate unified skew-elliptical $\operatorname{(SU\mathcal{E})}$ distribution, denoted by $\boldsymbol{X}\sim SU\mathcal{E}_{p,p}(\mu,\Sigma,\Lambda,h^{(p)},0,\Gamma)$, if its density function at $x\in\mathbb{R}^{p}$ is:
$f_{X}(x)=2^{p}f_{p}\left(x;\mu,\Sigma,h^{(p)}\right)F_{p}\left(\Lambda\sigma^{-1}(x-\mu);\Gamma,h_{Q(z)}^{(p)}\right),$ (5)
where $f_{p}\left(x;\mu,\Sigma,h^{(p)}\right)$ is the elliptical distribution defined in (4), and $F_{p}\left(\Lambda\sigma^{-1}(x-\mu);\Gamma,h^{(p)}_{Q(x)}\right)$ the $p$-dimensional centered elliptical cumulative distribution function with $p\times p$ dispersion matrix $\Gamma$, $\Lambda$ is a $p\times p$ real matrix controlling shape. In conclusion, $Q(\mathbf{x})$ is a quadratic form defined as $Q(\mathbf{x})=(\mathbf{x}-\mu)^{\top}\boldsymbol{\Sigma}^{-1}(\mathbf{x}-\mu)\in\mathbb{R}^{+}$, whereas $h_{Q(\mathbf{x})}^{(p)}(u)=h^{(2p)}[u+Q(\mathbf{x})]/g^{(p)}[Q(\mathbf{x})]$ denotes the elliptical conditional density generator.
@@
(arellano2010skew-e) Given $\boldsymbol{X}\sim\operatorname{SU\mathcal{E}}_{p,p}(\mu,\Sigma,\Lambda,h^{(p)},0,\Gamma)$, we can write $\boldsymbol{X}$ as a mixture of a CSN distribution: $\boldsymbol{X}=\mu+\sigma V_{0}^{-1/2}\boldsymbol{Z}_{0}$, where $V_{0}$ is the cumulative distribution function of a uni-dimensional random variable, called mixture variable, such that: $V_{0}(0)=0$, and $V_{0}$ is independent of $\boldsymbol{Z}_{0}\sim\operatorname{CSN}_{p,p}(0,\Sigma^{0},\Lambda,0,\Gamma)$.
@@
(arellano2010skew-e) If  $\text{ $V_{0}$}\sim Gamma(\nu/2,\nu/2)$, then the unified skew-elliptical distribution $\boldsymbol{X}\sim\operatorname{SU}\mathcal{E}_{p,p}(\mu,\Sigma,\Lambda,h^{(p)},0,\Gamma)$ reduces to a unified skew-$t$ $\operatorname{(SUT)}$ distribution: $\boldsymbol{X}\sim\operatorname{SUT}_{p,p}(\mu,\Sigma,\Lambda,\nu,0,\Gamma)$.
We construct an undirected graph on the closed skew-normal distribution, following zareifard2016skew.
@@
Consider an undirected graph $G=(V,E)$, a random vector $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$ is called a skew graphical model concerning the graph G with mean $\mu$, the precision matrix $\Omega$ and the skewness parameter $\alpha\in\mathbb{R}^{p}$ if its density is $\operatorname{CSN}_{p,p}(0,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$ or equivalently
@@
where $L$ and $D_{\kappa}$ are the matrix resulting from the modified Cholesky of $\Omega=LD_{\kappa}L^{\top}$, and zero in $L$ matrix determines the conditional independence relations between the $X_{1},\ldots,X_{p}$ variables. Further, when the matrix $D_{\alpha}=0$, the random vector $\boldsymbol{X}\sim\operatorname{CSN}_{p,p}(0,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$ reduces to $\boldsymbol{X}\sim\operatorname{N}_{p}(0,\Omega^{-1})$.
The graphical model for the unified skew-elliptical distribution can be constructed by utilizing the sparsity structure of the closed skew-normal distribution. We employ a semiparametric approach, utilizing the class of unified skew-elliptical distributions, to extend the SKEPTIC estimator proposed by liu2012high. This approach enables the development of robust estimators for graphical models that quantify the skewness present in the data distributions.
@@ 3 The elliptical skew-(S)KEPTIC estimators
(Meta skew-elliptical distribution) A continuous random vector $\boldsymbol{X}=(X_{1},\ldots,X_{p})^{\top}$ has meta skew-elliptical distribution, denoted by $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Sigma^{0},\Lambda,h^{(p)},0,\Gamma;f)$, if there exists a set of monotone univariate functions $f_{1},\ldots,f_{p}$ such that
$(f_{1}(X_{1}),\ldots,f_{p}(X_{p}))^{\top}\sim\operatorname{SU\mathcal{E}}_{p,p}(0,\Sigma^{0},\Lambda,h^{(p)},0,\Gamma),$ (7)
@@
where: $\boldsymbol{U}\sim HN_{p}(0,I_{p})$ and $\boldsymbol{V}\sim N_{p}(0,I_{p})$, where $HN_{p}(0,I_{p})$ is a $p$-dimensional half-normal distribution, and $N_{p}(0,I_{p})$ is a $p$-dimensional normal distribution, the distributions of $\boldsymbol{U}$ and $\boldsymbol{V}$ are independent. We introduce a correlation matrix estimator that explicitly accounts for the cross effects of univariate skewness parameters on the linear dependence among variables. The robustness of the proposed estimators follows from the direct estimation of the skewness parameters without the explicit computation of the marginal distributions, in the spirit of the SKEPTIC estimator of liu2012high.
@@
Assume that $\boldsymbol{X}$ follows a meta skew-elliptical distribution: $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Omega^{-1},D_{\alpha}L,h^{(p)},0,D_{\kappa};f)$, such that $\boldsymbol{Y}=f(\boldsymbol{X})$ denotes the unified skew-elliptical distribution $\boldsymbol{Y}\sim\operatorname{SU\mathcal{E}}_{p,p}(0,\Omega^{-1},D_{\alpha}L,h^{(p)},0,D_{\kappa})$. The elliptical skew-KEPTIC estimator is:
${S}^{\tau,\alpha}_{ij}=\begin{cases}\sin\Big(\frac{\pi}{2}\tau_{ij}\Big)\frac{1+\alpha_{i}\alpha_{j}}{\sqrt{(1+\alpha_{i}^{2})(1+\alpha_{j}^{2})}}&i\neq j\\ 1&i=j\end{cases}.$
As discussed in hult2002multivariate, Spearman’s rho is not invariant in the class of elliptical distributions, and its relation with the Pearson correlation coefficient in Lemma 2.1 holds if and only if the distribution is nonparanormal. The proof of the Theorem is in A.
@@
(elliptical skew-SKEPTIC estimator) Assume that $\boldsymbol{X}$ follows a meta skew-elliptical distribution: $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Omega^{-1},D_{\alpha}L,h^{(p)},0,D_{\kappa};f)$, such that $\boldsymbol{Y}=f(\boldsymbol{X})$ denotes the unified skew-elliptical distribution. When $\boldsymbol{Y}\sim\operatorname{SU\mathcal{E}}_{p,p}(0,\Omega^{-1},D_{\alpha}L,h^{(p)},0,D_{\kappa})$ reduces to the CSN distribution in (6), the elliptical skew-SKEPTIC estimator is:
${S}^{\rho,\alpha}_{ij}=\begin{cases}2\sin\Big(\frac{\pi}{6}\rho_{ij}\Big)\frac{1+\alpha_{i}\alpha_{j}}{\sqrt{(1+\alpha_{i}^{2})(1+\alpha_{j}^{2})}}&i\neq j\\ 1&i=j\end{cases},\quad{S}^{\tau,\alpha}_{ij}=\begin{cases}\sin\Big(\frac{\pi}{2}\tau_{ij}\Big)\frac{1+\alpha_{i}\alpha_{j}}{\sqrt{(1+\alpha_{i}^{2})(1+\alpha_{j}^{2})}}&i\neq j\\ 1&i=j\end{cases}.$
We generally refer to the estimators in Theorem 3.4 and Lemma 3.5 as elliptical skew-(S)KEPTIC estimators. By construction, the estimated dependence measures are amplified when the skewness components exhibit concordance, whereas they are attenuated when the skewness components are discordant. To obtain the correlation matrix using the elliptical skew-(S)KEPTIC estimators, we need to estimate the parameters of the distributions, which are encoded in the diagonal elements of $D_{\alpha}$, for closed skew-normal and unified skew-$t$ distributions.